# Efficient frontiers for portfolios under SSD and law-invariant risk measures with hyperbolic return distributions

### 05 Mar 2026 v9

@@ Abstract
The classical Markowitz mean–variance framework treats variance as the measure of risk and determines efficient frontier portfolios in closed form using standard optimization methods. For more general mean–risk settings, however, obtaining closed-form optimal portfolios is typically difficult. In this note, we derive explicit expressions for frontier portfolios corresponding to arbitrary law-invariant convex risk measures, assuming that the return vector follows a normal mean–variance mixture distribution. Our approach begins by developing stochastic dominance conditions within the family of normal mean–variance mixture models, which we then utilize to derive closed-form representations of frontier portfolios. The central result establishes that if asset returns are described by a normal mean–variance mixture model, then the associated mean–risk efficient frontier can be obtained by solving a Markowitz mean–variance problem with a suitably modified return vector. Moreover, we show that for any normal mean–variance mixture specification of the form $\eta=a+bZ+c\sqrt{Z}N$, the following relationship holds under every law-invariant coherent risk measure
@@
We then use this result to obtain a CAPM-style equilibrium model within our framework.
@@ 3 Frontier portfolios
$\hat{\rho}(X)=\sup_{\mu\in\mathcal{M}}\int_{(0,1]}CVaR_{\beta}(X)\mu(d\beta),$ (35)
@@ 4 Applications
By the positive homogeneity of $CVaR$, this yields the stated formula in the case of $CVaR$. For a general law-invariant convex risk measure $\rho$, the claim follows by applying relation (35). ∎
@@
The relation (36) above implies that, for any law-invariant convex risk measure $\rho$ (such as Expected Shortfall / CVaR), the risk of a mixture asset can be decomposed into two parts: the expectation of the linear term $\omega^{T}\mu+\omega^{T}\gamma EZ$ and the risk of the purely normal component, multiplied by the factor $\sqrt{\omega^{T}\Sigma\omega}$. This representation turns an otherwise intricate, distribution-dependent portfolio risk evaluation into a more tractable expression.
@@
We maximize the Sharpe ratio
$\frac{ER_{\omega}-r_{f}}{\rho(R_{\omega})}=\frac{\omega^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)}{\rho(R_{\omega})}.$
First note that, from (37) we have
$\frac{d\rho(R_{\omega})}{d\omega}=(\mu-\mathbf{1}r_{f}+\gamma EZ)+\frac{\rho_{0}}{2\sqrt{\omega^{T}\Sigma\omega}}\Sigma\omega.$
@@
for some function $L(\omega^{\star})$. Normalizing by the total weight, we obtain
@@
In what follows, we derive the corresponding capital asset pricing model (CAPM) based on the analysis developed in this paper. To do so, we begin by introducing the following risk measure:
for any NMVM random variable $\eta=a+b\eta+c\sqrt{Z}N$. Due to Proposition 4.2 we have $\bar{\sigma}(\eta)=c$.
@@
For any return $R=a+bZ+c\sqrt{Z}N$, we have
$ER-r_{f}=\frac{\rho(R)-ER}{\rho(R_{m})-ER_{m}}(ER_{m}-r_{f}).$ (39)
@@
We follow the idea of proof of the traditional CAPM. Consider a risky asset $i$ with return $R_{i}=a_{i}+b_{i}Z+c_{i}\sqrt{Z}N$ and the market return $R_{m}$. Invest a fraction $\omega$ of an investor’s wealth in asset $i$ and the remaining fraction $(1-\omega)$ in the market portfolio. The corresponding return is denoted by $R_{\omega}=\omega R_{i}+(1-\omega)R_{m}$. We have $ER_{\omega}=\omega ER_{i}+(1-\omega)ER_{M}$ and $\bar{\sigma}(R_{\omega})=\omega c_{i}+(1-\omega)\sigma_{m}$. We calculate $\frac{dER_{\omega}}{d\omega}=ER_{i}-ER_{m}$ and $\frac{d\bar{\sigma}}{d\omega}=c_{i}-\sigma_{m}$. We have $\frac{dER_{\omega}}{d\omega}|_{\omega=0}=ER_{i}-ER_{m}$ and $\frac{d\bar{\sigma}}{d\omega}|_{\omega=0}=c_{i}-\sigma_{m}$. We utilize the relation
$\frac{dER_{\omega}}{d\omega}|_{\omega=0}]/[\frac{d\bar{\sigma}}{d\omega}|_{\omega=0}]=\frac{ER_{m}-r_{f}}{\sigma_{m}}.$
This leads into the equation
$\frac{ER_{i}-ER_{m}}{c_{i}-\sigma_{m}}=\frac{ER_{m}-r_{f}}{\sigma_{m}}.$
After some algebra we obtain
$ER-r_{f}=\frac{c_{i}}{\sigma_{m}}(ER_{m}-r_{f}).$
Observe that, due to Proposition 4.2, we have $c_{i}=\frac{\rho(R_{i})-ER_{i}}{\rho_{0}}$ and $\sigma_{m}=\frac{\rho(R_{m})-ER_{m}}{\rho_{0}}$. This ends the proof. ∎
@@
Recall that the standard CAPM is
@@
in which the covariance between the asset return $R$ and the market return $R_{m}$ enters explicitly. By contrast, the skewness-induced CAPM (SI-CAPM) (39) does not feature this covariance term. Nonetheless, under our specification $R=a+bZ+c\sqrt{Z}N$, the return $R$ remains correlated with $R_{m}$, and in fact
$Cov(R_{m},R)=b\beta_{m}Var(Z)+c\sigma_{m}EZ.$
The classical CAPM relies on normally distributed returns and takes variance as the relevant notion of risk (which leads to beta as the risk measure). Our relation (39) yields an analogous linear link between expected return and risk. Even when returns follow a heavy-tailed mixture distribution (such as a t-distribution) and risk is assessed via a more intricate measure (such as Expected Shortfall), the resulting equilibrium pricing relation preserves the simple linear form
@@
However, in this framework, the beta $\beta$ is determined by the parameters of the mixture model rather than solely by the covariance with the market.
@@
The risk measure $\bar{\sigma}(\cdot)$ introduced in (38) is not convex. It can instead be interpreted as a risk measure analogous to the usual standard deviation. Under this risk measure, we clearly obtain $\bar{\sigma}(R_{\omega})=\sqrt{\omega^{T}\Sigma\omega}$ and $\bar{\sigma}^{2}(R_{\omega})=\omega^{T}\Sigma\omega$ (see also (37)). The tangent portfolio in Proposition 4.4 can likewise be derived using the risk measure $\bar{\sigma}^{2}(\cdot)$. The derivation proceeds in the same way as the construction of the tangent portfolio in the Markowitz setting, thereby providing additional justification for the computations carried out in Theorem 4.5.
Remark 4.8.
The formula (39) can be equivalently expressed as
$\frac{ER-r_{f}}{\rho(R)-ER}=\frac{ER_{m}-r_{f}}{\rho(R_{m})-ER_{m}},$
which shows that the ratio of the expected excess return $ER-r_{f}$ of any asset to its excess risk (over the expected return) $\rho(R)-ER$ is equal to the corresponding ratio for the market portfolio.
@@
$\mathcal{A}=(\mu+\gamma EZ)^{T}\Sigma^{-1}(\mu+\gamma EZ),\;\mathcal{B}=(\mu+\gamma EZ)^{T}\Sigma^{-1}\mathbf{1},\;\mathcal{C}=\mathbf{1}^{T}\Sigma^{-1}\mathbf{1},\;\triangle=\mathcal{A}\mathcal{C}-\mathcal{B}.$

### 16 Mar 2026 v10

@@ Abstract
The classical Markowitz mean–variance framework treats variance as the measure of risk and determines efficient frontier portfolios in closed form using standard optimization methods. For more general mean–risk settings, however, obtaining closed-form optimal portfolios is typically difficult. In this note, we derive explicit expressions for frontier portfolios corresponding to arbitrary law-invariant convex risk measures, assuming that the return vector follows a normal mean–variance mixture distribution. Our approach begins by developing stochastic dominance conditions within the family of normal mean–variance mixture models, which we then utilize to derive closed-form representations of frontier portfolios. The central result establishes that if asset returns are described by a normal mean–variance mixture model, then the associated mean–risk efficient frontier can be obtained by solving a Markowitz mean–variance problem with a suitably modified return vector. Furthermore, we demonstrate that for any normal mean–variance mixture of the form $\eta=a+bZ+c\sqrt{Z}N$, the following (unexpected) identity is valid for every law-invariant coherent risk measure:
@@
This relationship then serves as the basis for deriving a CAPM-type equilibrium model within our setting.
@@ 3 Frontier portfolios
$\hat{\rho}(X)=\sup_{\mathcal{P}\in\mathcal{M}}\int_{(0,1]}CVaR_{\beta}(X)\mathcal{P}(d\beta),$ (35)
@@ 4 Applications
By the positive homogeneity of $CVaR$, this yields the stated formula in the case of $CVaR$. For a general law-invariant coherent risk measure $\rho$, the claim follows by applying relation (35). ∎
@@
The relation (36) above implies that, for any law-invariant coherent risk measure $\rho$ (such as Expected Shortfall / CVaR), the risk of a mixture asset can be decomposed into two parts: the expectation of the linear term $\omega^{T}\mu+\omega^{T}\gamma EZ$ and the risk of the purely normal component, multiplied by the factor $\sqrt{\omega^{T}\Sigma\omega}$. This representation turns an otherwise intricate, distribution-dependent portfolio risk evaluation into a more tractable expression.
@@
We consider the maximization of the Sharpe ratio
$\frac{ER_{\omega}-r_{f}}{\rho(R_{\omega})}=\frac{\omega^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)}{\rho(R_{\omega})},$
over all portfolios satisfying $\omega^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)>0$. This condition guarantees that $\rho(R_{\omega})>0$. First, note that from (37) we have
$\frac{d\rho(R_{\omega})}{d\omega}=(\mu-\mathbf{1}r_{f}+\gamma EZ)+\frac{\rho_{0}}{\sqrt{\omega^{T}\Sigma\omega}}\Sigma\omega.$
@@
with
$L(\omega^{\star})=\frac{\rho(R_{\omega^{\star}})-(\omega^{\star})^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)}{(\omega^{\star})^{T}(\mu-\mathbf{1}r_{f}+\gamma EZ)}\frac{\sqrt{(\omega^{\star})^{T}\Sigma\omega^{\star}}}{\rho_{0}}.$
Normalizing by the total weight, we obtain
@@
In what follows, we examine several key consequences of the results established in this paper. We begin by defining the following risk measure:
for an arbitrary NMVM random variable $\eta=a+b\eta+c\sqrt{Z}N$. By Proposition 4.2, it follows that $\bar{\sigma}(\eta)=c$. The risk measure $\bar{\sigma}(\cdot)$ introduced in (38) is not convex. It can instead be interpreted as a risk measure analogous to the usual standard deviation. Under this risk measure, we clearly obtain $\bar{\sigma}(R_{\omega})=\sqrt{\omega^{T}\Sigma\omega}$ and $\bar{\sigma}^{2}(R_{\omega})=\omega^{T}\Sigma\omega$ (see also (37)). The tangent portfolio in Proposition 4.4 can likewise be derived using the risk measure $\bar{\sigma}^{2}(\cdot)$. The derivation proceeds in the same way as the construction of the tangent portfolio in the Markowitz setting, thereby providing additional justification for the computations carried out in Theorem 4.5 below.
@@
For any return of the form $R=a+bZ+c\sqrt{Z}N_{r}$, we have
$ER-r_{f}=\sigma_{rm}\frac{\rho(R)-ER}{\rho(R_{m})-ER_{m}}(ER_{m}-r_{f}),$ (39)
where $R_{m}=u_{m}+\beta_{m}Z+\sigma_{m}\sqrt{Z}N_{m}$ denotes the market return introduced in Proposition 4.4, and $\sigma_{rm}$ represents the correlation between the standard Normal random variable $N_{r}$ and the standard Normal variable $N_{m}$.
@@
We proceed along the lines of the standard CAPM proof. Consider a risky asset $i$ with return $R_{i}=a_{i}+b_{i}Z+c_{i}\sqrt{Z}N_{i}$ and a market portfolio with return $R_{m}=u_{m}+\beta_{m}Z+\sigma_{m}\sqrt{Z}N_{m}$. Let $\sigma_{im}$ be the correlation coefficient between the standard Normal variables $N_{i}$ and $N_{m}$. Suppose an investor allocates a fraction $\lambda$ of wealth to asset $i$ and the remaining fraction $(1-\lambda)$ to the market portfolio. The resulting portfolio return is $R_{\lambda}=\lambda R_{i}+(1-\lambda)R_{m}$. Its expected return is $ER_{\lambda}=\lambda ER_{i}+(1-\lambda)ER_{m},$ and its (modified) standard deviation is $\bar{\sigma}(R_{\lambda})=\sqrt{\lambda^{2}c_{i}^{2}+(1-\lambda)^{2}\sigma_{m}^{2}+2\lambda(1-\lambda)\sigma_{m}c_{i}\sigma_{im}}.$ We compute $\left.\frac{dER_{\lambda}}{d\lambda}\right|_{\lambda=0}=ER_{i}-ER_{m}$ and $\left.\frac{d\bar{\sigma}}{d\lambda}\right|_{\lambda=0}=\frac{\sigma_{m}c_{i}\sigma_{im}-\sigma_{m}^{2}}{\sigma_{m}}.$ Using the relation
$\left[\left.\frac{dER_{\lambda}}{d\lambda}\right|_{\lambda=0}\right]\Big/\left[\left.\frac{d\bar{\sigma}}{d\lambda}\right|_{\lambda=0}\right]=\frac{ER_{m}-r_{f}}{\sigma_{m}},$
we obtain
$ER_{i}-r_{f}=\frac{c_{i}\sigma_{im}}{\sigma_{m}}(ER_{m}-r_{f}).$
By Proposition 4.2, we know that
$c_{i}=\frac{\rho(R_{i})-ER_{i}}{\rho_{0}}\quad\text{and}\quad\sigma_{m}=\frac{\rho(R_{m})-ER_{m}}{\rho_{0}}.$
Substituting these expressions yields
$ER_{i}-r_{f}=\sigma_{im}\,\frac{\rho(R_{i})-ER_{i}}{\rho(R_{m})-ER_{m}}(ER_{m}-r_{f}).$ (40)
@@
Recall that the standard CAPM is given by
@@
where the covariance between the asset return $R$ and the market return $R_{m}$ appears explicitly. In contrast, the skewness-induced CAPM (SI-CAPM) (39) does not contain this covariance term. Instead, the SI-CAPM relies solely on the correlation between $N_{r}$ and $N_{m}$.
The traditional CAPM assumes normally distributed returns and uses variance as the key measure of risk, which leads to beta as the associated risk metric. Our relation (39) delivers a similar linear connection between expected return and risk. Even when returns follow a heavy-tailed mixture distribution (for instance, a t-distribution) and risk is measured by a more sophisticated criterion (such as Expected Shortfall), the equilibrium pricing relation still takes the simple linear form
@@
In this setting, however, the beta $\beta$ is determined by the mixture model parameters rather than being driven solely by the covariance with the market.
@@
If $N_{r}$ and $N_{m}$ in Theorem 4.5 are perfectly correlated, then $\sigma_{rm}=1$, and in this situation the expression (39) can be rewritten as
$\frac{ER-r_{f}}{\rho(R)-r_{f}}=\frac{ER_{m}-r_{f}}{\rho(R_{m})-r_{f}},$
indicating that the ratio of the expected excess return $ER-r_{f}$ of any asset to its excess risk $\rho(R)-r_{f}$ coincides with the corresponding ratio for the market portfolio. If $\sigma_{rm}<1$, then we obtain
$\frac{ER-r_{f}}{\rho(R)-ER}=\sigma_{rm}\frac{ER_{m}-r_{f}}{\rho(R_{m})-ER_{m}}.$
@@
$\mathcal{A}=(\mu+\gamma EZ)^{T}\Sigma^{-1}(\mu+\gamma EZ),\;\mathcal{B}=(\mu+\gamma EZ)^{T}\Sigma^{-1}\mathbf{1},\;\mathcal{C}=\mathbf{1}^{T}\Sigma^{-1}\mathbf{1},\;\triangle=\mathcal{A}\mathcal{C}-\mathcal{B}^{2}.$
@@
[…]
$\bar{\sigma}=\sqrt{\mathcal{C}r^{2}-2\mathcal{B}r+\mathcal{A}}.$
According to Theorem 3.4, this curve is obtained from mean–variance optimization using the return vector $\theta$ with mean $\mu_{\theta}=\mu+\gamma EZ$ and covariance matrix $\Sigma_{\theta}=\Sigma$. The curve is a hyperbola, and we obtain the straightforward relation
$\rho(r)=r+\frac{\rho_{0}}{\sqrt{\triangle}}\bar{\sigma}(r).$
In line with the findings of [2], the utility-maximizing optimal portfolios are located on this frontier.